# Investigating the Electroweak Phase Transition with a Real Scalar Singlet at a Muon Collider

### v1

@@ Abstract
A strong first-order electroweak phase transition (SFOEWPT) is essential for explaining baryogenesis and for potentially generating observable gravitational waves. This study investigates the potential of a high-energy muon collider to examine the occurrence of SFOEWPT within the context of a Standard Model extended by a real scalar singlet (xSM). We analyzed all possible decay modes of the singlet to constrain the valid parameter space of SFOEWPT, which was extracted numerically at different renormalization scales to account for theoretical uncertainties, thereby determining the sensitivity of a muon collider to the production and decay channels of novel heavy scalar particles that emerge in the xSM. The findings demonstrate that a 3 TeV muon collider can directly examine the nature of electroweak symmetry breaking by efficiently detecting novel scalar particles associated with a first-order electroweak phase transition through jet-rich final states, thus complementing the indirect constraints from gravitational wave experiments.
Keywords: Electroweak phase transition, real singlet extension, finite temperature correction, effective potential, Muon collider
@@ 1 Introduction
The current structure of the universe remains an unresolved enigma, tracing its origins to the initial moments of its existence, when symmetry breaking resulted in the differentiation of the fundamental forces observed today. Electroweak symmetry breaking (EWSB) occurred approximately $10^{-11}$ s after Big Bang, leading to the generation of mass for most Standard Model (SM) particles, followed by chiral symmetry breaking ($\sim 10^{-6}~{}\mathrm{s}$), which led to hadron confinement. Electroweak symmetry was spontaneously broken when the thermal evolution of the universe reached a state where the Higgs field’s potential became unstable around its vanishing minimum, leading to a transition towards another non-vanishing, “broken”, minimum that evolved at lower temperatures, a phenomenon referred to as the electroweak phase transition (EWPT) kirzhnits1972 ; coleman1973radiative ; dolan1974symmetry ; quiros1998 ; jackiw1974functional ; weinberg1973 ; dolan1974gauge . This transition could serve as Sakharov’s third condition for explaining baryogenesis, that is the generation of the observed matter-antimatter asymmetry if it was of the first-order type. Additionally, CP-violating processes that constitute Sakharov’s second condition would have to occur in thermal disequilibrium, in order to prevent the sphaleron wash-out of the generated asymmetric baryon number (Sakharov’s first condition) Sakharov:1967dj ; dine2003origin ; morrissey2012electroweak ; bhattacharyya2011pedagogical ; mclerran1991baryon ; senaha2013overview ; cline2006baryogenesis . Previous studies have shown that this sphaleron condition, which is quantitatively expressed as $\frac{E_{\text{\tiny sp}}(T_{c})}{T_{c}}>45$, can be translated into a relation between the theoretical parameters, $\frac{v(T_{c})}{T_{c}}\geq 1$, that can only be satisfied if a first-order EWPT (FOEWPT) occurs braibant1993sphalerons ; ahriche2007criterion . A FOEWPT is characterized by the existence of two degenerate minima that are separated by a barrier at a critical temperature $T_{c}$. In SM, such behavior can only arise from the cubic term in the Higgs field, generated from thermal loop corrections, which is further suppressed by a mass-screening effect resulting from the IR-divergences from the massless Matsubara modes. Earlier studies have suggested that SM can still produce FOEWPT if thermal corrections with the leading contribution of the IR-divergent daisy diagrams are included carrington1992effective . However, subsequent studies demonstrated that this possibility disappears when the subleading terms of the daisy and superdaisy contributions are included, which shifts the SM phase transition more towards the second-order type espinosa1993nature ; arnold1993effective . Currently, it is established that a large Higgs boson mass cannot allow for FOEWPT, therefore , at best, the SM EWPT can only be of crossover type when all non-perturbative corrections are included kajantie1996there ; senaha2020 . Thus, to explain spontaneous symmetry breaking (SSB) through FOEWPT, additional degree(s) of freedom must be added to the SM to improve the sphaleron condition, primarily by strengthening the Higgs field’s vacuum expectation value through coupling with other scalars.
@@
Multiple extensions have been proposed in recent decades, including additional real scalars, complex scalars, Higgs doublets, and supersymmetric extensions espinosa2012strong ; choi1993real ; o2007minimal ; ham2005electroweak ; ramsey2020electroweak ; papaefstathiou2022electro ; branco1998electroweak ; barger2009complex ; ELLWANGER20101 . The real gauge-singlet extension (xSM) is considered to be one of the simplest yet promising extensions to the Standard Model (SM) for several reasons. Primarily, it can enhance the critical cubic term through a direct, non-vanishing contribution at the tree level. Additionally, it serves as a simplified model for more complex extensions, such as that of a complex singlet or the Next-to-Minimal Supersymmetric Standard Model (NMSSM), in which multiple scalar fields exist with a substantial mass hierarchy. This hierarchy effectively isolates the heavier scalars from the electroweak symmetry breaking (EWSB) dynamics, with a contribution that is exponentially suppressed, resulting in only the lighter scalar being the effective scalar, which is adequately approximated by xSM ELLWANGER20101 . Conversely, this scenario presents significant testing challenges, often referred to as the “nightmare scenario” which renders it valuable for evaluating future colliders and assessing their capacity to make definitive statements regarding the nature of electroweak symmetry breaking. Consequently, a critical examination of xSM can effectively serve as a concurrent investigation of some of more complex extensions. For these reasons, this model has been extensively explored in previous research, including studies of its implications for the gravitational waves produced by real vacuum bubbles dynamics and its phenomenological consequences on the precision of the Higgs portal couplings at colliders ramsey2020electroweak ; carena2020electroweak ; papaefstathiou2022electro ; no2014probing ; huber2008gravitational ; curtin2014testing .
@@
While the nature of electroweak symmetry breaking is currently being explored at the Large Hadron Collider (LHC), it is also advantageous to examine its prospects in future collider facilities ramsey2020electroweak . A high-energy muon collider can provide sufficient energy and integrated luminosity, to obtain information complementary to that of an LHC al2022muon ; liu2021probing ; forslund2022high ; liu2021probing . In particular, the clean experimental environment of the muon collider will facilitate high-precision measurements in the Higgs sector. These measurements will not only test the SM but also provide information about potential extensions of the SM. This information should enable researchers to elucidate the mechanism of electroweak symmetry breaking and explore beyond the electroweak sector. Specifically, the maximal sensitivity of a lepton collider with center-of-mass energy $\sqrt{s}$, to a scalar of mass $m_{\phi}$, produced via vector boson fusion, occurs at approximately $\sqrt{s}/m_{\phi}\approx 1.7$ ramsey2020electroweak . Such a relationship does not exist for Hadron Colliders. Consequently, it is possible to optimize the search for a scalar in a muon collider by knowing its approximate mass. In the context of electroweak symmetry breaking, the mass of the scalar augmenting the strength of the phase transition is typically close to the electroweak scale, which is within the operational range of the muon collider. This characteristic renders the muon collider an ideal environment for discovering a new scalar associated with electroweak symmetry breaking. In this article, we present an analysis of the minimum sensitivity of a future muon collider to probe a real scalar singlet that could mediate a strong FOEWPT. We focus on the kinematical region where the new scalar mass exceeds that of the SM Higgs boson mass, thereby allowing for singlet decay into two Higgs bosons, consistent with the current LHC constraints. Subsequently, by analyzing the decay channels, we identified the most promising avenues for investigation in future muon colliders.
@@
The remainder of this paper is structured as follows: Section 2 aims to elucidate how the real singlet extension enhances the electroweak phase transition (EWPT) to be of a first-order type. Approximations were employed where appropriate to simplify calculations and highlight the primary features for pedagogical purposes. In addition, this section discusses the constraints utilized to identify the parameter space points that are consistent with FOEWPT. Section 3 addresses the direct and indirect verifications of xSM at colliders. Section 4 presents analysis of the various channels of the new heavy scalar decay. Finally, Section 5 discusses the results and provides commentary on the subsequent steps of this investigation.
@@ 2 Electroweak Phase Transition with a Scalar Singlet
$V_{int}=\frac{1}{2}\alpha S(H^{\dagger}H)+\frac{1}{4}\lambda_{hs}S^{2}(H^{\dagger}H)\;.$ (2)
A $\mathbb{Z}_{2}$-symmeteric potential can be obtained directly by setting the parameters of the odd term to zero ($\alpha,\beta\rightarrow 0$). In this case, if the mass of the resulting new scalar exceeds half of the Higgs boson mass, it becomes challenging to detect at the colliders. This phenomenon is commonly referred to as the ”nightmare scenario.” Nevertheless, in this study, we opted to keep these $\mathbb{Z}_{2}$-asymmeteric terms to maintain a more general representation that encompasses other types of challenging scenarios that induce some hard-to-reach regions of the parameter space, where both the mixing angle and di-Higgs boson decays are minimal. Furthermore, these specific terms are crucial for facilitating a strong first-order EWPT; see, for example choi1993real . The singlet term in the scalar field can be eliminated using shift symmetry, which only results in the redefinition of the remaining parameters, as demonstrated in previous studies  espinosa2012strong ; profumo2007singlet . Consequently, the most general renormalizable representation of the real gauge-singlet extension of SM (xSM) is,
@@
Electroweak symmetry breaking generates vacuum expectation values (vevs), $v$ and $\omega$, for the Higgs doublet and scalar singlet fields, respectively. The physical states are obtained by expanding around vevs, $H\rightarrow\frac{1}{\sqrt{2}}(v+h)$ and $S\rightarrow\omega+s$, derived from the minimization conditions:
@@
These equations can generate up to eight stationary points, in addition to the typical symmetric point at the origin($0,0$). Some of these points may be degenerate, whereas others may not be physically viable. Consequently, the singlet extension in Eq. (3) directly influences the Higgs vacuum at the tree level, which becomes dependent on the singlet parameters after spontaneous symmetry breaking (SSB):
@@
If the singlet has a vanishing vacuum expectation value (vev), $\omega=0$, or when $\alpha=\frac{1}{2}\lambda_{hs}\omega$, the broken minima in (7) will converge to the electroweak vev ,$v_{\text{\tiny EW}}=247$ GeV, which is also evident from Eq.(5). Moreover, Eq. (5) indicates that at high temperatures, where the vevs become temperature-dependent, a large negative value for the cubic term, $\alpha>\frac{1}{2}\lambda_{hs}\omega$, can derive the electroweak phase transition (EWPT) towards the first-order type solely through the tree-level (TL) contribution by improving the sphaleron condition. This is a consequence of a decrease in $\mu_{h}^{2}$ can significantly reduce the critical temperature as $T_{c}\propto\mu_{h}$ as will be demonstrated subsequently. This phenomenon may not be immediately apparent from (7) as a reduction in $\mu_{h}^{2}$ may result in decreasing $v_{b}$, and consequently may diminish the first-order electroweak phase transition (FOEWPT) condition, $\frac{v_{c}}{T_{c}}$, in totality. However, this is not the case, as will be demonstrated. Consequently, this extensive parameter space facilitates the possibility of obtaining FOEWPT. Conversely, the TL corrections would reduce the mixing angle and the di-Higgs boson decays, potentially necessitating higher luminosity to examine the $\mathbb{Z}_{2}$-asymmetric scenario compared with the $\mathbb{Z}_{2}$-symmetric one curtin2014testing ; ramsey2020electroweak . In addition to tree-level correction, NLO corrections, both at zero temperature from the Coleman-Weinberg one-particle irreducible (1PI) contribution, and at finite temperature, could further catalyze the phase transition towards FOEWPT. The Coleman-Weinberg correction to the tree level potential is given by
@@
The finite temperature corrections are obtained from (refer to, e.g. dolan1974symmetry ; quiros1998 ; laine2017basics for a comprehensive review),
@@
where, ${x=\frac{p}{T},\,y=\frac{m}{T}}$, and the sign depends on the particle type (negative for bosons and positive for fermions). For non-vanishing particle masses, $y\neq 0$, the integral must be evaluated numerically. Nevertheless, high-temperature expansion (HTE) to order $\mathcal{O}(y^{6})$,
@@
could be a valid approximation up to $m\lesssim 3T$ as explained in Ref. laine2017basics (See Fig. 1). In this study, we set $m_{s}\in[200:1000]$, indicating that certain scalar mass values do not conform to this approximation. However, for these values, the thermal corrections can be disregarded because in this case, $m_{s}>3T$, the low-temperature expansion will be a valid approximation, superseding the HTE, which is exponentially suppressed. Consequently, this contribution can be disregarded for this pedagogical analysis, particularly given that in $\mathbb{Z}_{2}$-asymmetric xSM, the FOEWPT is predominantly driven by the TL odd terms. Therefore, the calculations can be modified according to the valid mass as follows:
  • For $m_{s}\lesssim 3T$, which represents the majority of valid conservative parameter-space points (see Fig. 2), the thermal corrections can be approximated by the HTE, such that the overall effective potential becomes,
@@
  • For $m_{s}>3T$, which represents the majority of liberal parameter-space points, the overall effective potential becomes,
@@
Notably the fractional term proportional to $m_{i}^{4}$ in Eq. (12) and $m_{t}^{4}$ in Eq. (13) cancels it out with the corresponding CW correction in Eq. (8), whereas the logarithmic part is reduced to $\log\left[\frac{e^{2\gamma_{\text{\tiny E}}}T^{2}}{\Lambda^{2}}\right]$, where $\frac{e^{2\gamma_{\text{\tiny E}}}T^{2}}{\Lambda^{2}}\approx\mathcal{O}(1)$, if the scale is considered equal to the temperature. For other scales, one can qualitatively estimate the relative weight of the CW corrections by examining the possible shift in the Higgs vacuum structure when considering CW contributions as
@@
This results in an increase in coefficients of $v^{2}$ in Eq. (5) and Eq. (6), which is smaller than $0.005$ when only including the effect of $t,\,Z,\,W^{\pm}$. This becomes even smaller when the effects of $H,S$ are added because they have an overall negative value for the bracket in Eq. (19). Consequently, this correction can maximally change Higgs vev by a factor smaller than $1\%$ when compared to the leading $v^{2}$-coefficients; $\lambda_{h}\approx 0.13$ and $\alpha$ which could be even larger than $\lambda_{h}$ for non-vanishing $\omega$, when $\alpha>\frac{1}{2}\lambda_{hs}\omega$. Therefore, for an approximate understanding of the thermal evolution of Higgs vev, one can disregard the CW correction to the effective potential as well.
@@
This corresponds to the shift $m_{1,2}^{3}\rightarrow M_{1,2}^{3}=\left(m_{1,2}^{2}(v,\omega)+\frac{\lambda}{4}T^{2}\right)^{\frac{3}{2}}$, where the gauge bosons do not lead to any imaginary parts. Therefore, at very high temperatures, $\lambda T^{2}\gg m^{2}_{1,2}$, the scalar cubic term is reduced to a pure temperature term, independent of the fields. This implies that the cubic term coefficient will be exclusive to the vector gauge boson contributions, in this approximation, and will not receive any further corrections in xSM. Considering the main corrections in Eq. (11) and Eq. (20), the effective potential becomes,
@@
$+\frac{1}{4}\lambda_{h}v^{4}+\frac{1}{2}D(T^{2}-T_{1}^{2})\omega^{2}+\frac{1}{3}\beta\omega^{3}+\frac{1}{4}\lambda_{s}\omega^{4}+\mathcal{O}\left(\frac{1}{T^{2}}\right)\;$ (21)
@@
$E$ $=\frac{1}{32\pi}\left(2g_{1}^{2}+(g_{1}^{2}+g_{2}^{2})^{\frac{3}{2}}\right)\;,$ (23)
@@
and $\mathcal{O}\left(\frac{1}{T^{2}}\right)$ represent corrections arising from the $m^{6}$-term in the HTE. The vevs then become temperature dependent,
@@
where ${\xi(T)=\frac{(2\alpha+\lambda_{hs}\omega)\omega}{2\lambda_{h}T^{2}}}$. At critical temperature in SM, the sphaleron condition in Eq.26 returns ${\frac{v_{c}}{T_{c}}=\frac{2E}{\lambda_{h}}\approx 0.15}$, where ${T_{c}^{\text{\tiny SM}}=\frac{\mu_{h}}{\sqrt{C\left(1-\frac{2E^{2}}{\lambda_{h}C}\right)}}}$ and $\xi(T)=0$. This means that the sphaleron condition in xSM is primarly improved by the $\xi(T)-$correction term, which can contribute significant negative values to Eq.26 for large negative values of the odd $\mathbb{Z}_{2}-$asymmetric term $\alpha H^{\dagger}HS$. An additional supporting correction arises from the decrease in the critical temperature due to the $\lambda_{hs}$ portal coupling correction to the $C$-term, as per Eq.(22) and the correction to the $\mu_{h}^{2}$ term given in Eq.(5). The critical temperature is determined from the degeneracy condition [^2] We only consider a one-step phase transition, assuming that the scalar singlet field evolved a non-vanishing vev, which remains the same at $T_{c}$. In principle, a two-step phase transition mainly accounts for a shift in the effective potential that can affect the gravitational-wave spectrum, while barely affecting the value of the critical temperature carena2020electroweak .,

### v2

@@ Abstract
A strong first-order electroweak phase transition (SFOEWPT) is essential for explaining baryogenesis, and for potentially generating observable gravitational waves. In the present study, we investigate the potential of a high-energy muon collider to examine the occurrence of SFOEWPT within the context of a Standard Model extended by a real scalar singlet (xSM). We present an analysis of all viable decay modes of a singlet-like scalar particle, in order to constrain the valid parameter space of SFOEWPT, which was extracted numerically at different renormalization scales to account for theoretical uncertainties. This allowed us to determine the sensitivity of a muon collider to the production and decay channels of new heavy scalar singlet-like particles that emerge in the xSM. Our findings demonstrate that a 3 TeV muon collider could directly probe the nature of electroweak symmetry breaking by efficiently detecting new scalar particles associated with a first-order electroweak phase transition through jet-rich final states, thus complementing the indirect constraints from gravitational wave experiments.
Keywords: Electroweak phase transition, real singlet extension
@@ 1 Introduction
Several aspects of the current structure of the universe remain enigmatic, tracing their origins to the initial moments of its existence, when symmetry breaking resulted in the differentiation of the fundamental forces observed today. Electroweak symmetry breaking (EWSB) occurred approximately $10^{-11}$ s after Big Bang, leading to the generation of masses for most Standard Model (SM) particles, followed by chiral symmetry breaking ($\sim 10^{-3}~{}\mathrm{s}$), which led to hadron confinement. Electroweak symmetry was spontaneously broken when the thermal evolution of the universe reached a state where the Higgs field’s potential became unstable around its vanishing minimum, at the origin, leading to a transition towards another non-vanishing, “broken”, minimum that evolved at lower temperatures, a phenomenon referred to as the electroweak phase transition (EWPT) kirzhnits1972 ; coleman1973radiative ; dolan1974symmetry ; quiros1998 ; jackiw1974functional ; weinberg1973 ; dolan1974gauge . This transition could play the role of Sakharov’s third condition for explaining baryogenesis, i.e. the generation of the observed matter-antimatter asymmetry, if it was of the first-order type. Additionally, CP-violating processes that constitute Sakharov’s second condition would have to occur in thermal disequilibrium, in order to prevent the sphaleron wash-out of the generated asymmetric baryon number (Sakharov’s first condition) Sakharov:1967dj ; dine2003origin ; morrissey2012electroweak ; bhattacharyya2011pedagogical ; mclerran1991baryon ; senaha2013overview ; cline2006baryogenesis . Previous studies have shown that this sphaleron condition, which is quantitatively expressed as $\frac{E_{\text{\tiny sp}}(T_{c})}{T_{c}}>45$, can be translated into a relation between the theoretical parameters: $\frac{v(T_{c})}{T_{c}}\geq 1$. This condition can only be satisfied if a first-order EWPT (FOEWPT) occurs braibant1993sphalerons ; ahriche2007criterion . A FOEWPT is characterized by the existence of two degenerate minima that are separated by a barrier at some critical temperature $T_{c}$. In the SM, such behavior can only arise from the cubic term in the Higgs field, generated from thermal loop corrections, which is further suppressed by a mass-screening effect resulting from the IR-divergences from the massless Matsubara modes. Earlier studies have suggested that the SM can still produce a FOEWPT if thermal corrections with the leading contribution of the IR-divergent daisy diagrams are included carrington1992effective . However, subsequent studies demonstrated that this possibility disappears when the subleading terms of the daisy and superdaisy contributions are included, which shifts the SM phase transition more towards the second-order type espinosa1993nature ; arnold1993effective . Currently, it is established that a large Higgs boson mass cannot allow for FOEWPT, therefore, at best, the SM EWPT can only be of crossover type when all non-perturbative corrections are included kajantie1996there ; senaha2020 . Thus, to explain spontaneous symmetry breaking through a FOEWPT, additional degree(s) of freedom must be added to the SM to improve the sphaleron condition, primarily by strengthening the Higgs field’s vacuum expectation value through coupling with other scalar degrees of freedom.
@@
Multiple extensions have been proposed in recent decades, including additional real scalars, complex scalars, Higgs doublets, and supersymmetric extensions espinosa2012strong ; choi1993real ; o2007minimal ; ham2005electroweak ; ramsey2020electroweak ; papaefstathiou2022electro ; branco1998electroweak ; barger2009complex ; ELLWANGER20101 . The real gauge-singlet extension (xSM) is considered to be one of the simplest, yet promising, extensions to the Standard Model (SM) for several reasons. Primarily, it can enhance the sphaleron condition through direct correction to Higgs vacuum expectation value (vev) at the tree level. Additionally, it serves as a simplified model for more complex extensions, such as that of a complex singlet, or the Next-to-Minimal Supersymmetric Standard Model (NMSSM), in which multiple scalar fields exist with a substantial mass hierarchy. This hierarchy effectively isolates the heavier scalars from the electroweak symmetry breaking dynamics, with a contribution that is exponentially suppressed, resulting in only the lighter scalar being the effective scalar particle, which is adequately approximated by xSM ELLWANGER20101 . Conversely, this scenario presents significant testing challenges, often referred to as the “nightmare scenario”, which renders it valuable for evaluating future colliders and assessing their capacity to make definitive statements regarding the nature of electroweak symmetry breaking. Consequently, a critical examination of xSM can effectively serve as a concurrent investigation of some of more complex extensions. For these reasons, this model has been extensively explored in previous works, including studies of its implications for the gravitational waves produced by real vacuum bubble dynamics, and its phenomenological consequences on the precision of the Higgs portal couplings at colliders ramsey2020electroweak ; carena2020electroweak ; papaefstathiou2022electro ; no2014probing ; huber2008gravitational ; curtin2014testing .
@@
While the nature of electroweak symmetry breaking is currently being explored at the Large Hadron Collider (LHC), it is also advantageous to examine its prospects in future collider facilities ramsey2020electroweak . A high-energy muon collider can provide sufficient energy and integrated luminosity, to obtain information complementary to that of the LHC al2022muon ; liu2021probing ; forslund2022high . In particular, the clean experimental environment of the muon collider will facilitate high-precision measurements in the Higgs sector. These measurements can potentially provide valuable information regarding the possible extensions of SM, which may enable us to elucidate the mechanism of electroweak symmetry breaking. Furthermore, the direct detection of an additional scalar at a muon collider will facilitate more comprehensive analyses of all potential decay modes due to the enhanced reconstruction capabilities of final state events, including those with invisible final states. Consequently, it is possible to optimize the search for a scalar in a muon collider by knowing its approximate mass. In the context of electroweak symmetry breaking, the mass of the scalar augmenting the strength of the phase transition is typically close to the electroweak scale, which is within the operational range of a muon collider. This characteristic renders the muon collider an ideal environment for discovering a new scalar particle associated with electroweak symmetry breaking. In this article, we present an analysis of the minimum sensitivity of a future muon collider to probe a real scalar singlet field that could mediate a strong FOEWPT. We focus on the kinematical region where the new scalar particle’s mass exceeds that of the SM Higgs boson mass, thereby allowing for decays into two Higgs bosons, consistent with current LHC constraints. Subsequently, by analyzing the decay channels, we identify the most promising avenues for investigation in a future muon collider.
@@
The remainder of this paper is structured as follows: Section 2 aims to elucidate how the real singlet extension enhances the electroweak phase transition (EWPT) to be of a first-order type. Approximations were employed where appropriate to simplify calculations and highlight the primary features for pedagogical purposes. In addition, this section discusses the constraints utilized to identify the parameter space points that are consistent with FOEWPT. Section 3 addresses the direct and indirect verifications of xSM at colliders. Section 4 presents an analysis of the various channels of the new heavy scalar decay. Finally, Section 5 discusses the results and provides commentary on the subsequent steps of this investigation.
@@ 2 Electroweak Phase Transition with a Scalar Singlet
$V_{\mathrm{int}}=\frac{1}{2}\alpha S(H^{\dagger}H)+\frac{1}{4}\lambda_{hs}S^{2}(H^{\dagger}H)\;.$ (2)
A $\mathbb{Z}_{2}$-symmetric potential can be obtained directly by setting the parameters of the odd term to zero (i.e. $\alpha,\beta\rightarrow 0$). In this case, if the mass of the resulting new scalar exceeds half of the Higgs boson mass, it becomes challenging to detect at the colliders. This phenomenon is commonly referred to as the “nightmare scenario”. Nevertheless, in this study, we opted to keep these $\mathbb{Z}_{2}$-asymmeteric terms to maintain a more general representation that encompasses other types of challenging scenarios that induce some hard-to-reach regions of the parameter space, where both the mixing angle and di-Higgs boson decays are minimal. In addition, these specific terms are crucial for facilitating a strong first-order EWPT, see e.g. choi1993real . The singlet term in the scalar field can be eliminated using shift symmetry, which only results in the redefinition of the remaining parameters, as demonstrated in previous studies  espinosa2012strong ; profumo2007singlet . Consequently, the most general renormalizable representation of the real gauge-singlet extension of SM (xSM) is,
@@
Electroweak symmetry breaking generates vacuum expectation values (vevs), $v$ and $\omega$, for the Higgs doublet and scalar singlet fields, respectively. The physical states are obtained by expanding around the vevs, $H\rightarrow\frac{1}{\sqrt{2}}(v+h)$ and $S\rightarrow\omega+s$, derived from the minimization conditions:
@@
These equations can generate up to eight stationary points, in addition to the typical symmetric point at the origin, ($0,0$). Some of these points may be degenerate, whereas others may not be physically viable. Consequently, the singlet extension in Eq. (3) directly influences the Higgs vacuum at the tree level, which becomes dependent on the singlet parameters after spontaneous symmetry breaking:
@@
If the singlet has a vanishing vev, $\omega=0$, or when |$\absolutevalue{\alpha}=\frac{1}{2}\lambda_{hs}\omega$, the broken minima in Eq. (7) will converge to the electroweak vev, $v_{\text{\tiny EW}}\approx 246$ GeV, which is also evident from Eq.(5). Moreover, Eq. (5) indicates that at high temperatures, where the vevs become temperature-dependent, a large negative value for the cubic term, |$\absolutevalue{\alpha}>\frac{1}{2}\lambda_{hs}\omega$, can drive the electroweak phase transition towards the first-order type solely through the tree-level (TL) contribution by improving the sphaleron condition. This is a consequence of a decrease in $\mu_{h}^{2}$, and can significantly reduce the critical temperature as $T_{c}\propto\mu_{h}$ as will be demonstrated subsequently. This phenomenon may not be immediately apparent from Eq. (7) as a reduction in $\mu_{h}^{2}$ may decrease $v_{b}$, and consequently may diminish the first-order electroweak phase transition condition, $\frac{v_{c}}{T_{c}}$, in totality. However, this is not the case, as will be demonstrated. Consequently, this extensive parameter space facilitates the possibility of obtaining a FOEWPT. Conversely, the tree-level corrections would reduce the mixing angle and the di-Higgs boson decays for some parameter space points, potentially necessitating higher luminosity to examine the $\mathbb{Z}_{2}$-asymmetric scenario compared with the $\mathbb{Z}_{2}$-symmetric one curtin2014testing ; ramsey2020electroweak . In addition to tree-level corrections, NLO corrections, both at zero temperature from the Coleman-Weinberg one-particle irreducible (1PI) contribution, and at finite temperature, could further catalyze the phase transition towards a FOEWPT. The Coleman-Weinberg correction to the tree-level potential is given by
@@
The finite-temperature corrections are obtained from (refer to, e.g. dolan1974symmetry ; quiros1998 ; laine2017basics for a comprehensive review),
@@
where, ${x=\frac{p}{T},\,y=\frac{m}{T}}$, and the sign depends on the particle type (negative for bosons and positive for fermions). For non-vanishing particle masses, $y\neq 0$, the integral must be evaluated numerically. Nevertheless, a high-temperature expansion (HTE) to order $\mathcal{O}(y^{6})$,
@@
could be a valid approximation up to $m\lesssim 3T$ as explained in Ref. laine2017basics (See Fig. 1). In this study, we set $m_{s}\in[200:1000]$ GeV, indicating that certain scalar mass values do not conform to this approximation. However, for these values, the thermal corrections can be disregarded because in this case, $m_{s}>3T$, and the low-temperature expansion will be a valid approximation, superseding the HTE, which is exponentially suppressed. Consequently, this contribution can be disregarded for this pedagogical analysis, particularly given that in $\mathbb{Z}_{2}$-asymmetric xSM, the FOEWPT is predominantly driven by the tree-level odd terms. Therefore, the calculations can be modified according to the valid mass as follows:
  • For $m_{s}\lesssim 3T$, which represents the majority of valid conservative parameter space points (see Fig. 2), the thermal corrections can be approximated by the HTE, such that the overall effective potential becomes,
@@
  • For $m_{s}>3T$, which represents the majority of liberal parameter space points, the overall effective potential becomes,
@@
Notably, the fractional term proportional to $m_{i}^{4}$ in Eq. (12), and $m_{t}^{4}$ in Eq. (13), cancels out with the corresponding CW correction in Eq. (8), whereas the logarithmic part is reduced to $\log\left[\frac{e^{2\gamma_{\text{\tiny E}}}T^{2}}{\Lambda^{2}}\right]$, where $\frac{e^{2\gamma_{\text{\tiny E}}}T^{2}}{\Lambda^{2}}\approx\mathcal{O}(1)$, if the scale is considered equal to the temperature. For other scales, one can qualitatively estimate the relative weight of the CW corrections by examining the possible shift in the Higgs vacuum structure when considering CW contributions as
@@
This results in an increase in the coefficients of $v^{2}$ in Eq. (5) and Eq. (6), which is smaller than $0.005$ when only including the effect of $t,\,Z,\,W^{\pm}$. This becomes even smaller when the effects of $H,S$ are added because they have an overall negative value for the bracket in Eq. (19). Consequently, this correction can maximally change Higgs vev by a factor smaller than $1\%$ when compared to the leading $v^{2}$-coefficients; $\lambda_{h}\approx 0.13$ and $\alpha$ which could be even larger than $\lambda_{h}$ for non-vanishing $\omega$, when |$\absolutevalue{\alpha}>\frac{1}{2}\lambda_{hs}\omega$. Therefore, for an approximate understanding of the thermal evolution of Higgs vev, one can disregard the CW correction to the effective potential as well.
@@
This corresponds to the shift $m_{1,2}^{3}\rightarrow M_{1,2}^{3}=\left(m_{1,2}^{2}(v,\omega)+\frac{\lambda}{4}T^{2}\right)^{\frac{3}{2}}$, where the gauge bosons do not lead to any imaginary parts. Therefore, at very high temperatures, $\lambda T^{2}\gg m^{2}_{1,2}$, the scalar cubic term is reduced to a pure temperature term, independent of the fields. This implies that the cubic term coefficient will be exclusive to the vector gauge boson contributions in this approximation, and will not receive any further corrections in xSM. Considering the main corrections in Eq. (11) and Eq. (20), the effective potential becomes,
@@
$+\frac{1}{4}\lambda_{h}v^{4}+\frac{1}{2}D(T^{2}-T_{1}^{2})\omega^{2}+\frac{1}{3}\beta\omega^{3}+\frac{1}{4}\lambda_{s}\omega^{4}+\mathcal{O}\left(\frac{1}{T^{2}}\right)\;,$ (21)
@@
$E$ $=\frac{1}{32\pi}\left(2g_{1}^{2}+(g_{1}^{2}+g_{2}^{2})\right)^{\frac{3}{2}}\;,$ (23)
@@
and $\mathcal{O}\left(\frac{1}{T^{2}}\right)$ represents corrections arising from the $m^{6}$-term in the HTE. The vevs then become temperature dependent,
@@
where ${\xi(T)=\frac{(2\alpha+\lambda_{hs}\omega)\omega}{2\lambda_{h}T^{2}}}$. At the critical temperature in the SM, the sphaleron condition in Eq. 26 returns ${\frac{v_{c}}{T_{c}}=\frac{2E}{\lambda_{h}}\approx 0.15}$, where ${T_{c}^{\text{\tiny SM}}=\frac{\mu_{h}}{\sqrt{C\left(1-\frac{2E^{2}}{\lambda_{h}C}\right)}}}$ and $\xi(T)=0$. This means that the sphaleron condition in the xSM is primarly improved by the $\xi(T)-$correction term, which can contribute significant negative values to Eq. 26 for large negative values of the odd $\mathbb{Z}_{2}-$asymmetric term $\alpha H^{\dagger}HS$. An additional supporting correction arises from the decrease in the critical temperature due to the $\lambda_{hs}$ portal coupling correction to the $C$-term, as per Eq. (22) and the correction to the $\mu_{h}^{2}$ term given in Eq. (5). The critical temperature is determined from the degeneracy condition, [^2] We only consider a one-step phase transition, assuming that the scalar singlet field evolved a non-vanishing vev, which remains the same at $T_{c}$. In principle, a two-step phase transition mainly accounts for a shift in the effective potential that can affect the gravitational-wave spectrum, while barely affecting the value of the critical temperature carena2020electroweak .